# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v10

@@ 10.1 Network A/B Testing
The Seymour Second Neighborhood Conjecture is not just a theoretical problem; it has practical applications. We have constructed an algorithm that searches for and finds nodes that satisfy this conjecture. Not only does our algorithm find these nodes, but it partitions the graph into a data structure, which can be beneficial for many other purposes. While the conjecture only asked for a single node, our approach finds a set of degree-doubling nodes, which are critical in understanding both network topology and influence dynamics.
@@
The Graph Level Order (GLOVER) data structure is a way to organize an oriented graph into leveled neighborhoods. Nodes are placed based on their distance from a source node (normally a minimum out-degree node). Nodes are stored into ordered neighborhoods $R_{1}$, $R_{2}$, …, $R_{n}$, where $CR_{i}$ contains nodes that are distance $i$ from the root. Each neighborhood then connects to its next neighborhood.
The first thing this data structure does is helps with the problem of partitioning data sets. The data structure is defined with unique paths from $v_{0}$ to each hyper-node $R_{i}$. The hyper-nodes $R_{i}$ then connect only to the next hyper-nodes $R_{i+1}$. Thus, by the definition of the data structure, these sets $R_{i}$ and $R_{i+2}$ have no edges in common. This allows us to more easily set up the control group and the treatment group. The partitioning ensures that there can be no spill-over or network interference between the two groups. The GLOVER data structure, by organizing the graph based on distance from a source node, helps to minimize network interference. Because nodes in different neighborhoods (e.g., $R_{i}$ and $R_{i+2}$) have no direct connections, the likelihood of users in the treatment group influencing users in the control group is significantly reduced. This allows for more accurate measurement of the treatment’s true effect.
The GLOVER data structure, combined with our algorithm, allows for parallel processing of the graph. Because the neighborhoods in GLOVER are independent, the computation of degree-doubling nodes within each neighborhood can be performed concurrently. The ability to parallelize the algorithm significantly reduces the computation time, especially for very large social networks. This makes it possible to analyze the network and identify influential users in real-time or near real-time, which is essential for dynamic network A/B testing.
@@
Figure 19: Illustration showing partition of the GLOVER Data Structure into two groups, ready for Control and Treatment.
This application of the SSNC seeks to plant the seeds for future research in network A/B testing. There are other issues that the GLOVER algorithm does not resolve. Another limitation is that this algorithm will require an oriented graph, instead of a standard directed graph. that may limit applications as well. However, this research may hold a key to unlocking network A/B testing.
@@ 11 Conclusion
In this paper, we approached the Seymour Second Neighborhood Conjecture (SSNC) from a different, contradictory, angle. This allowed us to come up with a data structure that well ordered the nodes of the graph into neighborhoods. The data structure gave rise to a linear time algorithm that solves the SSNC and finds degree-doubling nodes. Most of the paper has focused on the mathematical aspects of this problem, trying to rigorously and constructively justify these claims. There are many real-world applications of both the GLOVER data structure and the algorithm. This was then shown to be applicable to network A/B testing, which makes the SSNC solution extremely relevant. Future work will look into investigating knowledge graphs for things like entities, emotions or hate speech in social media. This research is still ongoing and holds promise for understanding more about user behavior and particularly in the realm of natural language processing.
We have explored the Seymour Second Neighborhood Conjecture (SSNC) in oriented graphs. We did so through the lenses of graph theory, discrete mathematics, combinatorics, abstract algebra, algorithms, and data structures. This research helped us to use many mathematical tools, like both total orders, partitioning, divide and conquer, strong and weak induction, recurrence relations, and pathfinding techniques. One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph. This effectively solves the SSNC with an answer in the affirmative. The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree-doubling node was in the graph. Instead, the algorithm selects a minimum out-degree node and can partition the graph into neighborhoods and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
@@
Secondly, we introduced a data structure, Graph Level Order-GLOVER. This data structure allowed us to get a lot done on the conjecture. Many of the things inherent to SSNC made it easier for GLOVER—the well ordering and the dual metrics of distance and degrees. Still, the ability for the data structure to partition the nodes of an oriented graph into neighborhoods, which then allows for powerful methods like partitioning and mathematical induction, can have further applications. This would then allow us to implement resources from the field of graph theory into subject matter dealing with machine learning, data science, and network flow.
@@ 12 Acknowledgments
The GLOVER data structure is protected by a provisional patent application filed with the U.S. Patent and Trademark Office.

### v12

@@ 9.1 Network A/B Testing
The SSNC is not just a theoretical problem; it has practical applications. We have constructed an algorithm that searches for and finds nodes that satisfy this conjecture. Not only does our algorithm find these nodes, but it partitions the graph into a data structure, which can be beneficial for many other purposes. While the conjecture only asked for a single node, our approach finds a set of degree-doubling nodes, which are critical in understanding both network topology and influence dynamics.
@@
The Graph Level Order data structure is a way to organize an oriented graph into leveled neighborhoods. Nodes are placed based on their distance from a source node (normally a minimum out-degree node). Nodes are stored into ordered neighborhoods $R_{1}$, $R_{2}$, …, $R_{n}$, where $CR_{i}$ contains nodes that are distance $i$ from the root. Each neighborhood then connects to its next neighborhood.
The first thing this data structure does is helps with the problem of partitioning data sets. The data structure is defined with unique paths from $v_{0}$ to each hyper-node $R_{i}$. The hyper-nodes $R_{i}$ then connect only to the next hyper-nodes $R_{i+1}$. Thus, by the definition of the data structure, these sets $R_{i}$ and $R_{i+2}$ have no edges in common. This allows us to more easily set up the control group and the treatment group. The partitioning ensures that there can be no spill-over or network interference between the two groups. The Graph Level Order, by organizing the graph based on distance from a source node, helps to minimize network interference. Because nodes in different neighborhoods (e.g., $R_{i}$ and $R_{i+2}$) have no direct connections, the likelihood of users in the treatment group influencing users in the control group is significantly reduced. This allows for more accurate measurement of the treatment’s true effect.
The Graph Level Order, combined with our algorithm, allows for parallel processing of the graph. Because the neighborhoods in the Graph Level Order are independent, the computation of degree-doubling nodes within each neighborhood can be performed concurrently. The ability to parallelize the algorithm significantly reduces the computation time, especially for very large social networks. This makes it possible to analyze the network and identify influential users in real-time or near real-time, which is essential for dynamic network A/B testing.
@@
Figure 13: Illustration showing partition of the Graph Level Order into two groups, ready for Control and Treatment.
This application of the SSNC seeks to plant the seeds for future research in network A/B testing. There are other issues that the Graph Level Order algorithm does not resolve. One limitation is that this algorithm will require an oriented graph, instead of a standard directed graph. that may limit applications as well. However, this research may hold a key to unlocking network A/B testing.
@@ 10 Conclusion
In this paper, we approached the SSNC from a different, contradictory, angle. This allowed us to come up with a data structure that well ordered the nodes of the graph into neighborhoods. The data structure gave rise to a linear time algorithm that solves the SSNC and finds degree-doubling nodes. This data structure also adds to the mathematical literature on social media by dissecting transitive triangles into six distinct types that can help the SSNC with friendship recommendations. Most of the paper has focused on the mathematical aspects of this problem, trying to rigorously and constructively justify these claims. There are many real-world applications of both the Graph Level Order and the algorithm. This was then shown to be applicable to network A/B testing, which makes the SSNC solution extremely relevant. Future work will look into investigating knowledge graphs for things like entities, emotions or hate speech in social media. This research is still ongoing and holds promise for understanding more about user behavior and particularly in the realm of natural language processing.
This exploration of the SSNC has allowed us to look through the lenses of graph theory, discrete mathematics, combinatorics, abstract algebra, algorithms, and data structures. This research helped us to use many mathematical tools, like both total orders, partitioning, divide and conquer, strong and weak induction, recurrence relations, and pathfinding techniques. One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph. This effectively solves the SSNC with an answer in the affirmative. The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree-doubling node was in the graph. Instead, the algorithm selects a minimum out-degree node and can partition the graph into neighborhoods and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
@@
Secondly, we introduced a data structure, the Graph Level Order (GLOVER). This data structure allowed us to get a lot done on the conjecture. Many of the things inherent to SSNC made it easier for the Graph Level Order—the well ordering and the dual metrics of distance and degrees. Still, the ability for the data structure to partition the nodes of an oriented graph into neighborhoods, which then allows for powerful methods like partitioning and mathematical induction, can have further applications. This would then allow us to implement resources from the field of graph theory into subject matter dealing with machine learning, data science, and network flow.
@@ 11 Acknowledgments
I also appreciate the Department of Mathematics at Morehouse College for helpful discussions that were valuable during the research process. In particular, I thank Dr. Duane Cooper for asking the clarifying question where he asked for an example showing that Conjecture 1.1 and Conjecture 3.1 were equivalent.
@@
The Graph Level Order is protected by a provisional patent application filed with the U.S. Patent and Trademark Office.
Generative AI usage in the review process: Both ChatGPT 3.5 (July 20, 2024 version, published by OpenAI, https://chat.openai.com/) and Gemini (2.5 Flash version, Published by Google, https://gemini.google.com/) were used to analyze text drafts of human writing. All AI output was evaluated and judged based on merit by a human before being included anywhere in the manuscript.
Appendix A Appendix 1: Definitions
Definition A.1.
Definition A.2.
Definition A.3.
Definition A.4.
$N^{+}(v)=\{w\in V\mid(v,w)\in E\}$
Definition A.5.
Definition A.6.
Definition A.7.
Definition A.8.
A rooted neighborhood $R_{i}$ is the subgraph of $G$ induced by the set of nodes at distance $i$ from $v_{0}$, given a minimum out-degree node $v_{0}$. Formally,
$R_{i}=(V_{i},A_{i})$
where
$V_{i}=\{u\in V(G):\text{dist}(v_{0},u)=i\}$
and
$A_{i}=\{(u,v)\in E(G):u,v\in V_{i}\}$
where $V(G)$ and $E(G)$ denote the graph $G$’s vertex and edge sets, respectively, and $\text{dist}(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
Definition A.9.
Let $(x,y),(x,u),(y,u)\in G$. Then $x$, $y$, and $u$ form a transitive triangle, where $x$ is a common predecessor of both $y$ and $u$, and $y$ also connects to $u$.
Definition A.10.
Let $u_{i}$ be a node in the rooted neighborhood $R_{i}$ for some $i\geq 0$. A child of $u_{i}$ is a node $v_{i+1}\in R_{i+1}$ such that $(u_{i},v_{i+1})\in G$ (we also say that $u_{i}$ is the parent of $v_{i+1}$).
Definition A.11.
$\text{int}(u_{i},v_{1}):=N^{+}(u_{i})\cap N^{+}(v_{1})$
$deg_{\text{int}}(u_{i},v_{1}):=|\text{int}(u_{i},v_{1})|$
Definition A.12.
Let $u_{i}\in R_{i}$ be the parent of $v,w\in R_{i+1}$. Then $v$ and $w$ are said to be siblings.
Definition A.13.
Definition A.14.
Definition A.15.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
    • The interior neighbors of $u$ and $v$ are defined by the set $\text{int}(u,v)$.
    • The exterior neighbors of $u$ and $v$ are defined by the set $\text{ext}(u,v)$.
Definition A.16.
For a node $u\in G$ in an oriented graph $G$, we say that $u$ has the Decreasing Neighborhood Sequence Property if the size of its first out-neighbors is strictly larger than the size of its second out-neighbors, i.e., $|N^{++}(u)|<|N^{+}(u)|$.
Definition A.17.
Definition A.18.
Let $x\in R_{i}$. Define $\text{int}(R_{i-1},x)$ as the first out-neighbors of $x$ within $R_{i}$, and $\text{int}^{++}(R_{i-1},x)$ as the second out-neighbors of $x$ within $R_{i}$. We say that $x$ has its interior degree doubled if $|\text{int}^{++}(R_{i-1},x)|\geq|\text{int}(R_{i-1},x)|$.
Appendix B Appendix 2: Graph Level Order Applications
The deciding factor that held the algorithm together and helped solve the SSNC was the Graph Level Order data structure. This research did not begin as an investigation into data structures. Instead, a series of operations were conducted on the oriented graphs—partitioning, ordering those partitions, and adding interior and exterior arcs—until what was left was an ordering of both the partition and nodes within the partition. Further, these interior and exterior arcs allow us to define relationships based on two different metrics, an outer metric and an inner metric. Unlike the SSNC, exterior arcs do not always have to only be defined to the next neighborhood. That was a problem-specific definition.
Moving forward, the Graph Level Order should be well suited to tackle problems in graph theory and beyond. By introducing neighborhoods, another way of doing induction on graphs has been introduced. The concept of partitioning, which has dominated much of computer science, shows up in this paper as well. We can treat these neighborhoods independently of one another and solve their problems locally before bringing them back to the global problem. This shows how versatile the Graph Level Order is moving forward.
What makes Graph Level Order special are things like a total order on neighborhoods. Now there is a two-way ranking of nodes, as opposed to traditional one-way rankings like lexicographical sorting. This agrees with many real-world systems where we have dual metrics that are often competing, like price and performance. This is hard to measure on a single scale.
B.1 Representation
The more popular representations of graphs are adjacency matrices, array-lists, and edge lists. These suffer from the same limitations of graphs as mentioned above; they scatter the nodes along the two-dimensional plane without attempts to group them into an ordering. The Graph Level Order improves this by partitioning the nodes in a reasonable way. For the SSNC, that reasonable way was the distance metric. To represent graphs in a Graph Level Order, we can use any data serialization that can encode graphs, including JSON, YAML, and XML. There needs to be a new level of data: rooted neighborhoods, to go along with the standard data serialization. The neighborhoods will be given an ID associated with their distance from the minimum out-degree node. The nodes will then be assigned to their unique neighborhoods. Arcs will also be determined to be interior, exterior, or back, depending on the endpoints of the nodes of the arcs.
Node Targets Neighborhood
0 1, 2, 3 $R_{0}$
1 2, 4, 5, 6 $R_{1}$
2 3, 4, 5, 6 $R_{1}$
3 1, 4, 5, 6 $R_{1}$
4 5, 7, 8 $R_{2}$
5 6, 7, 8 $R_{2}$
6 7, 8, 1 $R_{3}$
7 8, 6, 1 $R_{3}$
8 6, 7, 1 $R_{3}$
Table 5: This figure illustrates a JSON representation of a Graph Level Order in array-list representations. Each node has a list of targets, along with its assigned neighborhood. Not shown are the nodes in neighborhood $R_{4}$. In this JSON format, each key represents a node ID, and its value is an object containing ’targets’ (a list of its out-neighbors) and ’neighborhood’ (the rooted neighborhood it belongs to, $R_{0}$ being the minimum out-degree node). This structure explicitly groups nodes by their distance from $v_{0}$.
With this rooted neighborhood representation of the SSNC, by way of the Graph Level Order, we can begin to visualize elements of these lemmas. For example, Lemma 6.8 (DNSP Neighborhood Size Property) can be visualized by Example 10. This shows a series of ordered rooted neighborhoods. Each rooted neighborhood has fewer nodes. The fewer nodes imply that the rooted neighborhood circle can be drawn with a smaller radius, giving light to the decrease in size.
Next, these nodes are not just treated like clusters inside these neighborhoods but like nodes in graph theory. The interior and exterior arcs allow for common graph theory algorithms and techniques like Prim’s algorithm, breadth-first search, or centrality to be called. What is even more interesting is that because we have identified the interior and exterior arcs already, we can choose to run these algorithms only on interior arcs or only on exterior arcs.
This is still a new data structure, so we are learning a lot about it. There are many applications that have not been tried yet. However, since it brings the strengths of a total order and a graph into one unified structure, it should be able to model complex relationships. This should include domains of machine learning like clustering, dynamic scheduling, resource management, project management, and natural language processing.
Figure 14: Here we have an illustration of some of the most common proofs in mathematics. They are partitioned by proof types. Interior arcs are drawn between nodes in the same neighborhood, representing things proven the same way. Exterior arcs are drawn between nodes that can be proven through multiple methods.
The Graph Level Order extends beyond oriented graphs though. This data structure can be used to represent arbitrary JSON/YAML/XML datasets. Anything that can be represented in one of these can be represented by a Graph Level Order and will have the tools of Graph Theory at hand.
Theorem B.1.
Any dataset that can be represented in a structured encoding format (such as JSON, XML, YAML, or equivalent) can be represented using a Graph Level Order, provided the user selects a valid metric. This will yield a valid Graph Level Order structure.
Proof.
Structured encoding formats, including JSON, XML, and YAML, encode data as discrete objects with finite fields and elements. Consequently, the number of elements in any dataset represented by these formats is finite, ensuring the data can be fully enumerated and partitioned.
Let $M$ be the metric chosen by the user such as degree or distance in the SSNC. $M$ is a field, attribute, or key present in the encoded data. $M$ corresponds to the basis for defining relationships across the neighborhoods.
For each element $x$ in the dataset, the metric $M(x)$ determines a value or set of values that can be used to group elements. The dataset is partitioned into neighborhoods $R_{1},R_{2},...,R_{k}$ based on $M$, where $R_{i}=\{x\mid M(x)=v_{i}\}$ for some value $v_{i}$.
If $M$ improperly defines neighborhoods (e.g., allows elements to belong to multiple neighborhoods or leaves elements unassigned), then $M$ is not a valid metric. A valid $M$ ensures that all elements are correctly assigned to exactly one neighborhood.
Thus, any dataset encoded in a structured format can be represented by a Graph Level Order, provided a suitable metric $M$ exists. ∎
The thing about the Graph Level Order is the adaptability. It adapts to the environment, and in particular the partition of that environment. SSNC called for a partition that needed an anchor node. In general most partitions do not require this and so we will not be seeing these types of graphs in the future unless especially called for. What we will be seeing are requests for partitions. Numeric partitions like distance, generally give rise to orderings easier. Other partitions may give rise to lexicographic orderings. However, the largest benefit of the Graph Level Order moving forward is extending graph theory concepts beyond standard graph theory and into the world of JSON/YAML/XML.
Appendix C Appendix 3: Visualization and Accessibility of Oriented Graphs
When we take a course on graph theory, oriented graphs are probably not on the course load. Second out-neighbors or squares of graphs are probably not either. To fully grasp this problem, though, we need to be able to delve into these definitions through visualizations.
For example, research on the SSNC led this paper into the terms Decreasing Neighborhood Sequence Property, induced subgraphs, neighborhoods, interior out-neighbors, exterior out-neighbors, and back arcs. The paper addressed how these terms were intertwined with each other. An example is where the paper says that there is a monotonically decreasing sequence.
James Robert Brown’s book ”Philosophy of Mathematics” speaks of the importance of visualizations in mathematics. Proofs are still the foundation, but sometimes visualizations can serve as inspiration or a spark for those proofs.
This realization motivated the creation of a complimentary website. The site serves as an interactive tool to understand the research. It also has things like MathJax installed to read lemmas, definitions, and proofs from this paper. Where possible, things like HTML canvas and JavaScript D3.js are used to visualize these concepts and help improve understanding.
Readers are encouraged to explore the website for additional information and examples at
https://glovermethod.com.
This is an evolving project. Additional examples and applications will continue to be added.
Readers are invited to explore the website. Please provide feedback on any aspects of the site or content for improvement.
Our goal here was to help users connect with the SSNC. Hopefully, this tool not only serves as support for this paper but also as an educational resource in the fields of graph theory and computer science. Thank you for your engagement.